The sum of three integers is
270. The sum of the first and second integers exceeds the third by
98.
The third integer is
47 less than the first.
Find the three integers

Answer by fractalier(6550)

Call the three integers x, y, and z. Then we write
x + y + z = 270
x + y = z + 98
z = x - 47
Now substitute the second into the first and get
z+98 + z = 270
2z + 98 = 270
2z = 172
z = 86
Then
86 = x - 47 and
x = 133
Plugging x and z in to the first gives us y...
133 + y + 86 = 270
219 + y = 270
y = 51
Your numbers are 133, 51 and 86.